# Vortical waves in a quantum fluid with vector, axial and helical charges. II. Dissipative effects

### v1

@@ 5.3 Dispersion relations
$DQ_{\ell}+Q_{\ell}\theta+\omega^{\mu}\partial_{\mu}\sigma^{\omega}_{\ell}+\sigma^{\omega}_{\ell}\partial_{\mu}\omega^{\mu}+\partial_{\mu}V^{\mu}_{\ell;d}=$ $0,$
@@
$(E+P)Du^{\mu}-\nabla^{\mu}P+\Delta^{\mu}{}_{\lambda}\partial_{\nu}\pi^{\lambda\nu}_{d}=$ $0,$ (148)
that can be cast with the help of the relations (146) into the following form in the rotating fluid:
$\pi^{\mu\nu}=$ $\frac{2i\eta k}{3}\begin{pmatrix}0&-y\Omega&x\Omega&0\\ -y\Omega&1&0&-\tfrac{3}{2}yv\Omega\\ x\Omega&0&1&\frac{3}{2}xv\Omega\\ 0&-\frac{3}{2}yv\Omega&\frac{3}{2}xv\Omega&-2\end{pmatrix}e^{-i(\omega t-kz)}\delta u,$
$V^{\mu}_{\ell}=$ $-i\tau_{R}\begin{pmatrix}0\\ y\Omega\omega\\ -x\Omega\omega\\ k\end{pmatrix}e^{-i(\omega t-kz)}\left(\frac{1}{3}\delta Q_{\ell}-\frac{Q_{\ell}\delta P}{E+P}\right).$ (149)
Furthermore, taking into account that $u_{\mu}\partial_{\nu}\pi^{\mu\nu}=-\pi^{\mu\nu}\nabla_{\nu}u_{\mu}$ is of second order in perturbations, while $\Delta^{\mu}{}_{\lambda}\partial_{\nu}\pi^{\lambda\nu}=\frac{4\eta k^{2}}{3}\delta u(0,0,0,1)^{\mu}e^{-i(\omega t-kz)}$, it is clear that the energy-momentum sector remains decoupled from the charge currents sector, since it involves only the amplitudes of perturbations in the velocity $\delta u$ and and the pressure $\delta P$:
$\begin{pmatrix}-3\omega&4Pk\\ k&-4P\omega-\frac{4}{3}i\eta k^{2}\end{pmatrix}\begin{pmatrix}\delta P\\ \delta u\end{pmatrix}=0.$ (150)
The above equation reveals that the effect of the dissipating shear is to shift the angular frequencies of the acoustic modes by an imaginary quantity, such that the non-dissipative solution of Eq. (13) becomes
$\omega^{\pm}_{R;{\rm ac.}}=\pm kc_{s}(\eta)-\frac{ik^{2}\eta}{6P},\qquad c_{s}(\eta)=\frac{1}{\sqrt{3}}\sqrt{1-\frac{k^{2}\eta^{2}}{12P^{2}}}.$ (151)
In other words, one of the effects of the presence of a finite kinetic relaxation time $\tau_{R}$ is to set the attenuation of the acoustic waves due to the nonvanishing kinetic shear viscosity (145). In addition, Eqs. (145) and (151) reveal that the speed of sound in the dissipative fluid is slower than the one in its non-dissipative limit (13). The speed of the sound wave depends on the momentum $k$, which has an ultraviolet bound, implying that when $k\tau_{R}>5\sqrt{3}/2$, the acoustic wave stops propagating. The latter effect is a feature of the first-order theory, as the inclusion of second-order terms may alter this behaviour Sammet:2023bfo .
What are the effects of a finite kinetic relaxation time for the propagation of perturbations in the charge sector? In the case of charge (non-)conservation, Eq. (12b) is modified to:
@@
$\left[\left(\omega+\frac{ik^{2}\tau_{R}}{3}\right)\delta_{\ell,\ell^{\prime}}-\Omega\frac{\partial\sigma^{\omega}_{\ell}}{\partial Q_{\ell^{\prime}}}\right]\delta Q_{\ell^{\prime}}-\left(\frac{ik^{2}\tau_{R}Q_{\ell}}{4P}+k\Omega\frac{\partial\sigma^{\omega}_{\ell}}{\partial P}\right)\delta P-(kQ_{\ell}-2\omega\Omega\sigma_{\ell}^{\omega})\delta u=-\frac{iT^{2}\delta\mu_{\ell}}{3\tau_{\ell}}.$ (152)
As in the case of the perfect fluid, the charge modes have vanishing pressure and velocity fluctuations, $\delta P=\delta u=0$, such that the last two terms above cancel. This leads to
@@
$\left[\left(\omega+\frac{ik^{2}\tau_{R}}{3}\right)\delta Q_{\ell}-k\Omega\delta\sigma^{\omega}_{\ell}\right]\biggl{|}_{\delta P=0}=-\frac{iT^{2}\delta\mu_{\ell}}{3\tau_{\ell}}.$ (153)
@@
where $\omega$ is the mode angular frequency.
@@ 5.4 Estimation of damping due to kinetic effects
Refer to caption Refer to caption
Figure 11: Perturbations in the chemical potentials at a fixed time $t=5$ fm/c with finite helical ($\tau_{H}=2.7$ fm/c) and axial ($\tau_{A}=0.25$ fm/c) relaxation times as we increase the angular frequency $\Omega$. The system is initialized with a Gaussian perturbation in the vector chemical potential $\delta\mu_{V}$ of width $\sigma=0.2$ fm on top of a neutral unpolarized background $\mu_{V}=\mu_{A}=\mu_{H}=0$ at temperature $T=300$ MeV.
@@
To illustrate the effect of the rotation parameter $\Omega$ on the propagation properties of the vortical waves, we considered a system initialized with a Gaussian perturbation in the vector chemical potential $\delta\mu_{V}$ of width $\sigma=0.2$ fm. Figure 11 shows the vector and helical perturbations at a time $t=5$ fm/c after initialization. We do not show the axial perturbation, as in the neutral plasma, it is not excited. It can be seen that at small angular frequency $\Omega$, the dissipative effects dominate the evolution of the perturbations while increasing $\Omega$ enhances the propagating features of the waves.
Refer to caption Refer to caption
Figure 12: Perturbations in the chemical potentials at a fixed time $t=5$ fm/c as we increase the damping due to kinetic effects, similar to the one of Fig. 11. We choose $\sigma=0.2$ fm for the width of initial Gaussian perturbations in all the chemical potentials $\delta\mu_{\ell}$ on top of a neutral background $\mu_{V}=0$ MeV at temperature $T=300\ {\rm M}_{\rm e}\!{\rm V}$ and the angular frequency $\Omega=100$ MeV for the finite helical ($\tau_{H}=2.7$ fm/c) and axial ($\tau_{A}=0.25$ fm/c) relaxation times.
We now estimate the effects of kinetic dissipation, governed by the kinetic relaxation time $\tau_{R}$. To estimate $\tau_{R}$, we considering that the ratio between the shear viscosity $\eta=\frac{4}{5}\tau_{R}P$ and the entropy density $s=(E+P-\vec{\mu}\cdot\vec{Q})/T$ is constant. According to Ref. Kovtun:2004de , the ratio $\eta/s$ should be bounded from below by $1/4\pi$ for strongly-interacting theories. Bayesian estimates based on matching hydrodynamics simulations to experimental data from heavy-ion collisions give $4\pi\eta/s\lesssim 1.5$ around $T=150$ MeV Bernhard:2019bmu , thus confirming this conjecture. Taking $14\pi\eta/s=$ 1 for definiteness and considering the choice of parameters discussed above ($T=300$ MeV, $\mu_{V}=0$), we obtain
@@
Considering the configuration and parameters shown in Fig. 11, we now vary $\eta/s$ while keeping $\Omega=100$ MeV fixed. The case $\tau_{R}=0$ is already depicted in Fig. 11. Figure 12 illustrate the effect of the kinetic dissipation as $\eta/s$ from 00 to $1/(4\pi)$. The figure confirms that kinetic dissipation damps short wavelength perturbations, smoothing the propagating peaks and progressively erasing the propagating features of the waves as we increase the $\eta/s$ ratio.
Before ending this section, we come back to our initial considerations and take into account realistic values $\Omega$. As pointed out in Ref. STAR:2017ckg , the angular velocity is estimated at $\Omega\simeq 6.6\ {\rm MeV}$ for the fireball created in non-central heavy-ion collisions. At this value of $\Omega$, the propagation speed of the helical vortical wave becomes a mere
$c_{h}{\biggl{|}}_{\Omega=6.6\ {\rm MeV}}=\frac{6\Omega\ln 2}{\pi^{2}T}\simeq 0.0093\,c.$ (158)
@@ 6 Summary and Conclusions
In our paper, we further developed an alternative approach to chiral fermionic systems, put forward in Refs. Ambrus:2019ayb ; Ambrus:2019khr ; HVWp1 , which extends the traditional premise that chiral fluids are fully described by the pair of vector and axial local charges. In the previous works, we have demonstrated that the incorporation of the helical degree of freedom of non-interacting relativistic fermions – which is a conserved degree of freedom related but not equal to the axial charge in general – enriches the hydrodynamic spectrum of the system via interplay for Axial and Helical Vortical Effects visualized in an intuitively appealing pictorial form in Fig. 1 of Ref. HVWp1 .
@@
In Ref. Ambrus:2019khr , we uncovered a new type of gapless mode, the Helical Vortical Wave, which represents a coherent propagating excitation of vector and helical charge densities and their currents. We also found the Axial Vortical Wave HVWp1 , which is driven by fluctuations in the axial charge density that propagates in the background of a large axial chemical potential $\mu_{A}$. Moreover, a background with non-zero $\mu_{A}$ induces non-reciprocity in the spectrum, namely that waves travel with different velocities along and opposite to the direction of the angular velocity. Contrary to the Chiral Vortical Wave Jiang:2015cva , which operates in vector and axial sectors, the Axial and Helical Vortical Waves emerge in an electrically neutral regime characterized by a vanishing vector chemical potential. In the opposite, high-density limit, these waves merge and form a mixed Axial-Vortical Wave HVWp1 . In this work, we question whether these hydrodynamic modes could withstand realistic conditions that appear in the interacting vortical quark-gluon plasmas.
@@
In our analysis, we used the relaxation time approximation in which interactions can phenomenologically be described in a relaxation time approximation via the introduction of the helical, $\tau_{H}$, and axial, $\tau_{A}$, and kinetic, $\tau_{R}$, relaxation times. The first two quantities control the time scale of the non-conservation of the corresponding charges, while the third one determines the dissipative evolution of occupation numbers out of the thermal equilibrium. The relaxation time of the vector charge is taken to be infinite as the electric charge is a conserved quantity. This property allowed us to study finite-density systems at non-zero vector chemical potential. Accordingly, in this paper, we do not consider the linear hydrodynamic waves at a finite background axial density since this case is not consistent with a finite axial relaxation time.
@@
We first analyzed the genuine effects of finite helical relaxation time $\tau_{H}$ in the limit when other relaxation effects are absent, $\tau_{A}=\tau_{R}=\infty$ (Section 3). In the dilute plasma regime, where the density of the vector chemical potential is small compared to temperature, $\mu_{V}\ll T$, and the other chemical potentials are zero, $\mu_{A}=\mu_{H}=0$, we get an expected effect: in addition to the damping due to the finite $\tau_{H}$, the Helical Vortical Wave evolves to a purely diffusive mode when the half-period of the wave exceeds the helical relaxation time $\tau_{H}$. In other words, the helical relaxation cuts off the low-frequency “helical” modes, corresponding predominantly to the helical-vortical charge fluctuations of the Helical Vortical Wave. The “axial” mode, a leftover of the non-reciprocal Axial Vortical Wave branch, is purely diffusive in this regime.
@@
The Chiral Vortical Wave emerges in the vanishing-$\tau_{H}$ limit where the helical degree of freedom disappears from the spectrum (gets frozen). It is worth stressing that the Chiral Vortical Wave and the mixed Axial-Helical Vortical Wave, both existing in the high-density limit at a finite $\tau_{H}$, are different excitations, as is reflected in their different propagation velocities and the chemical content of these waves.
Notice also that for a degenerate ($\mu_{V}\gg T$) axially conserved ($\tau_{A}\to\infty$) chiral fluid, the limit of the fast helical relaxation, $\tau_{H}\to 0$, is accompanied by the redistribution of the degrees of freedom. At large but finite $\tau_{H}$ (approximately conserved helicity), the spectrum is represented by two non-reciprocal modes of the Axial-Helical Wave that mixes almost indistinguishable helical-axial oscillations with fluctuations in the vector charge density. However, at small $\tau_{H}$, these modes get transformed to a non-propagating helical mode (dissipation of all charges driven by the helical charge relaxation) and two bi-directionally propagating axial-vector modes (the Chiral Vortical Wave).
At a finite (non-zero and non-infinite) helical relaxation time, the interplay of the axial and helical modes in dense fermionic matter becomes rather non-trivial, as summarized in Fig. 4. For the chemical potential below a certain limiting value, the wave spectrum contains three regions: (i) two propagating waves with a small wavenumber (a large wavelength) are the generalization of the Chiral Vortical Wave (axial-vector oscillations) that includes an accompanying helical mode; (ii) two propagating waves with a large wavenumber which is essentially a generalization of the Helical Vortical Wave (helical-vector oscillations) to the finite-density case; (iii) an intermediate region with a moderate wavelength in which both mentioned modes are diffusive. Thus, we have two branches of hydrodynamic excitations. When the vector chemical potential exceeds the mentioned critical value, the diffusive region (iii) disappears, and the (i) CVE and (ii) HVW domains merge, constituting a single branch of solutions. In all mentioned regions, these two helical modes are accompanied by a single diffusive axial non-propagating mode, a reminder of the Axial Vortical Wave.
@@
The accounting for a finite axial relaxation time $\mu_{A}$ makes the whole picture more realistic and, somewhat expectedly, more intricate. While we refer an interested reader to Section 5 for more details, here we describe the main features of the hydrodynamic waves with finite helical and axial relaxation times.
@@
If the helical charge dissipates rapidly (faster than the helical charge), then a single non-propagating helical mode dissipates rather quickly, while the two axial modes propagate at small wavevectors and are purely diffusive otherwise. In the case of conserved helicity, $\tau_{H}\to\infty$ (and at finite $\tau_{A}$), two non-dissipative, for all wavelengths, helical modes co-exist along with the purely dissipative axial mode.

### v2

@@ 5.3 Dispersion relations
$DQ_{\ell}+Q_{\ell}\theta+\partial_{\mu}(\omega^{\mu}\sigma^{\omega}_{\ell})+\partial_{\mu}V^{\mu}_{\ell;d}$ $=-\frac{T^{2}\mu_{\ell}}{3\tau_{\ell}},$
@@
$(E+P)Du^{\mu}-\nabla^{\mu}P+\Delta^{\mu}{}_{\lambda}\partial_{\nu}\pi^{\lambda\nu}_{d}$ $=0.$ (144)
We follow a similar approach as in Ref. Morales-Tejera:2024uzg and study the waves generated by longitudinal fluctuations of the four-velocity to leading order in $\Omega$. In particular, we write
$u^{\mu}=$ $\partial_{t}+\Omega(-y\partial_{x}+x\partial_{y})$
$+\Omega\delta u^{\perp}\left(-y\partial_{x}+x\partial_{y}\right)+\delta u^{z}\partial_{z}\,,$ (145)
where the $(x,y)$-dependence of the transverse plane fluctuations is required to have a consistent solution to leading order in $\Omega$, as shown in Ref. Morales-Tejera:2024uzg . The fluctuations are Fourier transformed in the $(t,z)$ coordinates as in Eq. (11). Then, Eqs. (144) in Fourier space can be cast with the help of the relations (142) into the following form in the rotating fluid:
$\pi_{d}^{\mu\nu}$ $\to i\eta k\delta u^{\perp}_{\omega}\begin{pmatrix}0&0&0&0\\ 0&0&0&y\Omega\\ 0&0&0&-x\Omega\\ 0&y\Omega&-x\Omega&0\end{pmatrix}$
$+\frac{2i\eta k}{3}\delta u^{z}_{\omega}\begin{pmatrix}0&-y\Omega&x\Omega&0\\ -y\Omega&1&0&-\tfrac{3\omega}{2k}y\Omega\\ x\Omega&0&1&\tfrac{3\omega}{2k}x\Omega\\ 0&-\tfrac{3\omega}{2k}y\Omega&\tfrac{3\omega}{2k}x\Omega&-2\end{pmatrix},$
$V^{\mu}_{\ell;d}$ $\to-i\tau_{R}\begin{pmatrix}0\\ y\Omega\omega\\ -x\Omega\omega\\ k\end{pmatrix}\left(\frac{1}{3}\delta Q_{\ell;\omega}-\frac{Q_{\ell}\delta P_{\omega}}{E+P}\right).$ (146)
Furthermore, taking into account that $u_{\mu}\partial_{\nu}\pi^{\mu\nu}=-\pi^{\mu\nu}\nabla_{\nu}u_{\mu}$ is of second order in perturbations, while
@@
$\Delta^{\mu}{}_{\lambda}\partial_{\nu}\pi^{\lambda\nu}\to\frac{4\eta k^{2}}{3}\delta u^{z}_{\omega}\delta^{\mu}_{z}\\ +\eta k\Omega\left(k\delta u^{\perp}_{\omega}-\frac{5}{3}\omega\delta u^{z}_{\omega}\right)(x\delta^{\mu}_{y}-y\delta^{\mu}_{x}).$ (147)
It is clear that the energy-momentum sector remains decoupled from the charge currents sector, since it involves only the amplitudes of perturbations in the velocity $\delta u^{\mu}_{\omega}$ and in the pressure $\delta P_{\omega}$. We now expand the angular frequency in terms of $\Omega$ as $\omega=\omega_{0}+\Omega\omega_{1}+\dots$, and similarly we decompose the fluctuations $\delta P_{\omega}$ and $\delta u^{z}_{\omega}$. To leading order in $\Omega$, the conservation equations for the energy momentum tensor reduce to
$\begin{pmatrix}-3\omega_{0}&4Pk\\ k&-4P\omega_{0}-\frac{4}{3}i\eta k^{2}\end{pmatrix}\begin{pmatrix}\delta P_{\omega;0}\\ \delta u^{z}_{\omega;0}\end{pmatrix}=0.$ (148)
Imposing a non-trivial solution to the previous equation reveals that
$\omega^{\pm}_{0}$ =±$=\pm kc_{s}(\eta)-\frac{ik^{2}\eta}{6P},$
$c_{s}(\eta)$ $=\frac{1}{\sqrt{3}}\sqrt{1-\frac{k^{2}\eta^{2}}{12P^{2}}},$ (149)
while the perturbations are related to each other as
$\delta P_{\pm;0}=\frac{4kP}{3\omega_{0}^{\pm}}\delta u^{z}_{\pm;0}.$ (150)
We have chosen $\delta u^{z}_{\omega;0}$ as our reference amplitude, and therefore we can set $\delta u^{z}_{\omega;1}=0$ without loss of generality. Then, the subleading contribution to the conservation equations for the energy momentum tensor take the following form:
$3\omega_{0}^{\pm}\delta P_{\pm;1}+\dfrac{4kP\omega_{1}^{\pm}}{\omega_{0}^{\pm}}\delta u^{z}_{\pm;0}$ $=0\,,$
$k\delta P_{\pm;1}-4P\omega_{1}^{\pm}\delta u^{z}_{\pm;0}$ $=0\,,$
$(12iP\omega_{0}^{\pm}-3k^{2}\eta)\delta u^{\perp}_{\pm}+(4iP+\eta\omega_{0}^{\pm})k\delta u^{z}_{\pm;0}$ $=0\,.$ (151)
The last equation can be used to solve for the transverse perturbation $\delta u^{\perp}_{\pm}$, while the first two equations indicate that the non-trivial solution ($\delta u^{z}_{\pm;0}\neq 0$) requires $\omega_{1}^{\pm}=0$. Consequently, the dispersion relation for the longitudinal sound mode does not get linear corrections in $\Omega$. The above discussion reveals that the effect of the dissipating shear is to shift the angular frequencies of the acoustic modes by an imaginary quantity, such that the non-dissipative solution shown in Eq. (12) becomes
$\omega^{\pm}_{R;{\rm ac.}}$ =±$=\pm kc_{s}(\eta)-\frac{ik^{2}\eta}{6P},$
$c_{s}(\eta)$ $=\frac{1}{\sqrt{3}}\sqrt{1-\frac{k^{2}\eta^{2}}{12P^{2}}}.$ (152)
The modification to the dispersion relation due to dissipation, $\omega^{\pm}_{R;{\rm ac.}}-\omega^{\pm}_{\rm ac.}=-ik^{2}\eta/6P=-i2k^{2}\tau_{R}/15$, agrees with that derived in Eq. (58) of Ref. Gorbar:2017toh using Chiral Kinetic Theory. [^5] A small discrepancy in the $O(\Omega)$ correction to $\omega^{\pm}_{R;{\rm ac.}}$, which can be traced back to the transverse plane dynamics ignored in Ref. Gorbar:2017toh , is irrelevant in the present case, as we work under the assumption of a chirally-neutral background. In other words, one of the effects of the presence of a finite kinetic relaxation time $\tau_{R}$ is to set the attenuation of the acoustic waves due to the nonvanishing kinetic shear viscosity (141). In addition, Eqs. (141) and (152) reveal that the speed of sound in the dissipative fluid is slower than the one in its non-dissipative limit (12). The speed of the sound wave depends on the momentum $k$, which has an ultraviolet bound, implying that when $k\tau_{R}>5\sqrt{3}/2$, the acoustic wave stops propagating. The latter effect is a feature of the first-order theory, as the inclusion of second-order terms may alter this behaviour Sammet:2023bfo .
What are the effects of a finite kinetic relaxation time for the propagation of perturbations in the charge sector? In the case of charge (non-)conservation, Eq. (13) is modified to:
@@
$\left[\left(\omega+\frac{ik^{2}\tau_{R}}{3}\right)\delta Q_{\ell}-k\Omega\delta\sigma^{\omega}_{\ell}\right]\biggl{\rvert}_{\delta P=0}=-\frac{iT^{2}\delta\mu_{\ell}}{3\tau_{\ell}}.$ (153)
@@
where $\omega$ is the mode angular frequency. The modification of the dispersion relation by the amount $ik^{2}\tau_{R}/3$ is consistent with the results reported in Eq. (59) of Ref. Gorbar:2017toh .
@@ 5.4 Estimation of damping due to kinetic effects
Figure 11: Perturbations in the chemical potentials at a fixed time $t=5$ fm/c with finite helical ($\tau_{H}=2.7$ fm/c) and axial ($\tau_{A}=0.25$ fm/c) relaxation times and vanishing $\tau_{R}$, as we increase the angular frequency $\Omega$. The system is initialized with a Gaussian perturbation in the vector chemical potential $\delta\mu_{V}$ of width $\sigma=0.2$ fm on top of a neutral unpolarized background $\mu_{V}=\mu_{A}=\mu_{H}=0$ at temperature $T=300$ MeV.
@@
To illustrate the effect of the rotation parameter $\Omega$ on the propagation properties of the vortical waves, we considered a system initialized with a Gaussian perturbation in the vector chemical potential $\delta\mu_{V}$ of width $\sigma=0.2$ fm. Figure 11 shows the vector and helical perturbations at a time $t=5$ fm/c after initialization. We do not show the axial perturbation, since in the neutral plasma, it is not excited. It can be seen that at small angular frequency $\Omega$, the dissipative effects dominate the evolution of the perturbations while increasing $\Omega$ enhances the propagating features of the waves.
Figure 12: Perturbations in the chemical potentials at a fixed time $t=5$ fm/c as we increase the damping due to kinetic effects, similar to the one of Fig. 11. We choose $\sigma=0.2$ fm for the width of initial Gaussian perturbations in all the chemical potentials $\delta\mu_{\ell}$ on top of a neutral background $\mu_{V}=0$ MeV at temperature $T=300$ MeV and the angular frequency $\Omega=100$ MeV for the finite helical ($\tau_{H}=2.7$ fm/c) and axial ($\tau_{A}=0.25$ fm/c) relaxation times.
We now estimate the effects of kinetic dissipation, governed by the kinetic relaxation time $\tau_{R}$. To estimate $\tau_{R}$, we consider that the ratio between the shear viscosity $\eta=\frac{4}{5}\tau_{R}P$ and the entropy density $s=(E+P-\vec{\mu}\cdot\vec{Q})/T$ is constant. According to Ref. Kovtun:2004de , the ratio $\eta/s$ should be bounded from below by $1/4\pi$ for strongly-interacting theories. Bayesian estimates based on matching hydrodynamics simulations to experimental data from heavy-ion collisions give $4\pi\eta/s\lesssim 1.5$ around $T=150$ MeV Bernhard:2019bmu , in agreement with this conjecture. Taking $14\pi\eta/s=$ 1 for definiteness and considering the choice of parameters discussed above ($T=300$ MeV, $\mu_{V}=0$), we obtain
@@
Considering the configuration and parameters shown in Fig. 11, we now vary $\eta/s$ while keeping $\Omega=100$ MeV fixed. The case $\tau_{R}=0$ is already depicted in Fig. 11. Figure 12 illustrates the effect of the kinetic dissipation as $\eta/s$ is increased from 00 to $1/(4\pi)$. The figure confirms that kinetic dissipation damps short wavelength perturbations, smoothing the propagating peaks and progressively erasing the propagating features of the waves as the $\eta/s$ ratio increases.
Before ending this section, we come back to our initial considerations and take into account realistic values of $\Omega$. As pointed out in Ref. STAR:2017ckg , the angular velocity is estimated at $\Omega\simeq 6.6\ {\rm MeV}$ for the fireball created in non-central heavy-ion collisions. At this value of $\Omega$, the propagation speed of the helical vortical wave becomes a mere
$c_{h}{\biggl{\rvert}}_{\Omega=6.6\ {\rm MeV}}=\frac{6\Omega\ln 2}{\pi^{2}T}\simeq 0.0093\,c.$ (158)
@@ 6 Summary and Conclusions
In our paper, we further developed an alternative approach to chiral fermionic systems, put forward in Refs. Ambrus:2019ayb ; Ambrus:2019khr ; Morales-Tejera:2024uzg , which extends the traditional premise that chiral fluids are fully described by the pair of vector and axial local charges. In the previous works, we have demonstrated that the incorporation of the helical degree of freedom of non-interacting relativistic fermions – which is a conserved degree of freedom related but not equal to the axial charge in general – enriches the hydrodynamic spectrum of the system via the interplay of Axial and Helical Vortical Effects, visualized in an intuitively appealing pictorial form in Fig. 1 of Ref. Morales-Tejera:2024uzg .
@@
In Ref. Ambrus:2019khr , we uncovered a new type of gapless mode, the Helical Vortical Wave, which represents a coherent propagating excitation of vector and helical charge densities and their currents. We also found the Axial Vortical Wave Morales-Tejera:2024uzg , which is driven by fluctuations in the axial charge density that propagates in the background of a large axial chemical potential $\mu_{A}$. Moreover, a background with non-zero $\mu_{A}$ induces non-reciprocity in the spectrum, namely that waves travel with different velocities along and opposite to the direction of the angular velocity. Contrary to the Chiral Vortical Wave Jiang:2015cva ; Gorbar:2017toh , which operates in vector and axial sectors, the Axial and Helical Vortical Waves emerge in an electrically neutral regime characterized by a vanishing vector chemical potential. In the opposite, high-density limit, these waves merge and form a mixed Axial-Helical-Vortical Wave Morales-Tejera:2024uzg . In this work, we questioned whether these hydrodynamic modes could withstand realistic conditions that appear in the interacting vortical quark-gluon plasmas.
@@
In our analysis, we used the relaxation time approximation in which interactions can phenomenologically be described via the introduction of the helical ($\tau_{H}$), axial ($\tau_{A}$) and kinetic ($\tau_{R}$) relaxation times. The first two quantities control the time scale of the non-conservation of the corresponding charges, while the third one determines the dissipative evolution of the fluid towards thermal equilibrium. The relaxation time of the vector charge is taken to be infinite as the electric charge is a conserved quantity. This property allowed us to study finite-density systems at non-zero vector chemical potential. Accordingly, in this paper, we do not consider the linear hydrodynamic waves at a finite background axial density since this case is not consistent with a finite axial relaxation time.
@@
We first analyzed the genuine effects of finite helical relaxation time $\tau_{H}$ in the limit when other relaxation effects are absent, i.e. $\tau_{A}=\infty$ and $\tau_{R}=0$ (Section 3). In the dilute plasma regime, where the density of the vector chemical potential is small compared to temperature, $\mu_{V}\ll T$, and the other chemical potentials are zero, $\mu_{A}=\mu_{H}=0$, we get an expected effect: in addition to the damping due to the finite $\tau_{H}$, the Helical Vortical Wave evolves to a purely diffusive mode when the half-period of the wave exceeds the helical relaxation time $\tau_{H}$. In other words, the helical relaxation cuts off the low-frequency “helical” modes, corresponding predominantly to the helical-vortical charge fluctuations of the Helical Vortical Wave. The “axial” mode, a leftover of the non-reciprocal Axial Vortical Wave branch, is purely diffusive in this regime.
@@
The Chiral Vortical Wave emerges in the vanishing-$\tau_{H}$ limit where the helical degree of freedom disappears from the spectrum (it freezes). It is worth stressing that the Chiral Vortical Wave and the mixed Axial-Helical Vortical Wave, both existing in the high-density limit at finite $\tau_{H}$, are different excitations, as is reflected in their different propagation velocities and the chemical content of these waves.
Notice also that for a degenerate ($\mu_{V}\gg T$) axially conserved ($\tau_{A}\to\infty$) chiral fluid, the limit of the fast helical relaxation, $\tau_{H}\to 0$, is accompanied by the redistribution of the degrees of freedom. At large but finite $\tau_{H}$ (approximately conserved helicity), the spectrum is represented by two non-reciprocal modes of the Axial-Helical Wave that mixes almost indistinguishable helical-axial oscillations with fluctuations in the vector charge density. However, at small $\tau_{H}$, these modes transform into a non-propagating helical mode (dissipation of all charges driven by the helical charge relaxation) and two bi-directionally propagating axial-vector modes (the Chiral Vortical Wave).
At a finite (non-zero and non-infinite) helical relaxation time, the interplay of the axial and helical modes in dense fermionic matter becomes rather non-trivial, as summarized in Fig. 4. For the chemical potential below a certain limiting value, the wave spectrum contains three regions: (i) two propagating waves with a small wavenumber (large wavelength) are the generalization of the Chiral Vortical Wave (axial-vector oscillations) that includes an accompanying helical mode; (ii) two propagating waves with a large wavenumber which is essentially a generalization of the Helical Vortical Wave (helical-vector oscillations) to the finite-density case; (iii) an intermediate region with a moderate wavelength in which both mentioned modes are diffusive. Thus, we have two branches of hydrodynamic excitations. (iii) When the vector chemical potential exceeds the mentioned critical value, the diffusive region disappears, and the (i) CVW and (ii) HVW domains merge, constituting a single branch of solutions. In all mentioned regions, these two helical modes are accompanied by a single diffusive axial non-propagating mode, a remainder of the Axial Vortical Wave.
@@
The accounting for a finite axial relaxation time $\mu_{A}$ makes the whole picture more realistic and, somewhat expectedly, more intricate. While we refer the interested reader to Section 5 for more details, here we describe the main features of the hydrodynamic waves with finite helical and axial relaxation times.
@@
If the helical charge dissipates rapidly (faster than the axial charge), then a single non-propagating helical mode dissipates rather quickly, while the two axial modes propagate at small wavevectors and are purely diffusive otherwise. In the case of conserved helicity, $\tau_{H}\to\infty$ (and at finite $\tau_{A}$), two non-dissipative helical modes co-exist along with the purely dissipative axial mode, for all wavelengths.